Dividing Partnership Income Tyler Hawes and Piper Albright formed a partnership, investing $210,000 and $70,000, respectively. Determine their participation in the year’s net income of $290,000 under each of the following independent assumptions: No agreement concerning division of net income. Divided in the ratio of original capital investment. Interest at the rate of 5% allowed on original investments and the remainder divided in the ratio of 2:3. Salary allowances of $36,000 and $45,000, respectively, and the balance divided equally. Allowance of interest at the rate of 5% on original investments, salary allowances of $36,000 and $45,000, respectively, and the remainder divided equally. Hawes Albright (a) $ 145,000 $ 145,000 (b) $ 217,500 $ 72,500 (c) $ $ (d) $ $ (e) $ $
Answer #1

(a) $ 145,000 $ 145,000

Divide $290000 equally to both partners

(b) $ 217,500 $ 72,500

As per investment Ratio will be 290000 X 210000/280000 and 290000 X 70000/280000

(c) $120900 $169100

Interest at the rate of 5% allowed on original investments and the remainder divided in the ratio of 2:3.

Interest payable is Hawes 210000 X 5/100 = 10500 and Albright 70000 X 5/100 = 3500

Remaining net income will be 290000-10500-3500=276000

Dividing 276000 in the ratio 2:3 will be

Hawes 276000 X 2/5 =110400

Albright 276000 X 3/5 = 165600

Hawes share will be 10500+110400=120900

Albright share will be 3500+165600=169100                     

(d) $140500 $149500

Salary allowances of $36,000 and $45,000, respectively, and the balance divided equally.

Remaining income will be 290000-36000-45000=209000 i.e both will get 104500

Hawes share will be 36000(salary)+104500=140500

Albright share will be 45000(salary)+104500=149500

(e) $144000 $146000

Allowance of interest at the rate of 5% on original investments, salary allowances of $36,000 and $45,000, respectively, and the remainder divided equally

Interest payable is Hawes 210000 X 5/100 = 10500 and Albright 70000 X 5/100 = 3500

Salary allowances of $36,000 and $45,000

Remaining income will be 290000-10500-3500-36000-45000=195000

195000 divided equally will be 97500

Hawes share will be 36000(salary)+10500 (interest) + 97500(Income share)=144000

Albright share will be 45000(salary)+3500 (interest) + 97500(Income share)=146000